OpenStudy (anonymous):

)Dawn is making an abstract painting with two triangles. The dimensions of the painting are shown below: Two triangles share the same height of 8 inches. One triangle has a base of 6 inches and the other is 10 inches. The total area of the two triangles is _____ square inches.

OpenStudy (anonymous):

I got 480 inch I want someone to check my answer :3 I will fan and medal

OpenStudy (anonymous):

@anna_truong123 @quickstudent

OpenStudy (anonymous):

@Empty @TheGaMer007 @wio @Ibbutibbu.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

You can do this 2 ways. 1. Calculate the area of each triangle and add them together. \(A = \dfrac{bh}{2} \) \(A = \dfrac{6 \times 8}{2} + \dfrac{10 \times 8}{2} = 24 + 40 = 64 in.^2\)

OpenStudy (anonymous):

yep the answer will be 64

OpenStudy (mathstudent55):

2. Calculate the area of the big triangle. \(A = \dfrac {bh}{2} = \dfrac{(6 + 10)8}{2} = \dfrac{16 \times 8}{2} = 64 ~in.^2\)\)

OpenStudy (anonymous):

Ok thank you